Quotients Homophones des Groupes Libres Homophonic Quotients of Free Groups

Jean-François Mestre René Schoof Lawrence Washington Don Zagier

Abstract

Soit G le quotient du groupe libre à 26 générateurs a, b, c, …, z par les relations A = B, pour tout couple de mots (A, B) pouvant avoir la même prononciation en anglais. Notre but est de déterminer la structure du groupe G. Notre résultat principal est le suivant:

Théorème 1A. G est trivial.

Voir aussi le théorème 1F ci-dessous.

Let G be the quotient of the free group on 26 letters a, b, c, …, z by the relations A = B whenever A and B are words having the same pronunciation in French. The object of this paper is to determine the structure of G. The main result is this:

Theorem 1F. G is trivial.

See also Theorem 1A above.